Find u xv, v xu, and v x v. v = (-5, 4,6) U = (9, -3, -2), (a) U XV (b) VXU (c) VXV
Answer #1

Solutions U=(9,-3,-2), Val-5,4,6) 1 i uXV j 9 -3 4 6 uxva (=18+8) -3654-10) +8536-15) uxus -loi-44+211 UXU=(-10,-44,21). ® VX